OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (anonymous):

hope it helps f(g(2))=-7

OpenStudy (solomonzelman):

When you are asked this type of questions. For example. \(\large\color{blue}{ \bf f(x)=2x+3 }\) and \(\large\color{blue}{ \bf g(x)=x-1 }\) \(\large\color{blue}{ \bf f(g(x)) = 2(x-1)+3 }\) \(\large\color{blue}{ \bf f(g(x)) = 2x-2+3 }\) \(\large\color{blue}{ \bf f(g(x)) = 2x-1 }\) Now, \(\large\color{blue}{ \bf f(g(4)) = 2(4)-1 }\) \(\large\color{blue}{ \bf f(g(4)) = 8-1 }\) \(\large\color{blue}{ \bf f(g(4)) =7 }\)